[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jun 08, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jun 8 18:45:42 UTC 2017


HSU Shift Log for Thu Jun 08 2017 14:13:27 GMT-0400 (EDT)

Shift: H

Forecaster: L Avila

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: A hot line test was successfully performed at 18z.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170608/76d31679/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list